oregonduck (oregonduck):

HELP WITH FLVS ALGEBRA 1

oregonduck (oregonduck):

Part B: If at one time the profit of the company was at least $60, what domain could possibly produce this profit? (2 points) Part C: What is an approximate average rate of change of the graph from x = 3 to x = 5, and what does this rate represent? (3 points)

OpenStudy (anonymous):

do u take it? just asking

oregonduck (oregonduck):

yes

OpenStudy (anonymous):

i saw a while back you had an argument with @bbaby and you said you were in college? LOL

oregonduck (oregonduck):

i know i was lying i was just trying to get at him

oregonduck (oregonduck):

can you help me?

OpenStudy (anonymous):

oh ok LMAO

OpenStudy (anonymous):

wait one sec brb

oregonduck (oregonduck):

ok

oregonduck (oregonduck):

@aouefbaowbu

oregonduck (oregonduck):

brb

OpenStudy (ccswims):

what test is this?

OpenStudy (ccswims):

what module?

oregonduck (oregonduck):

module 9 module test part 2

OpenStudy (anonymous):

Part B: When y is 60 x is both 1 and 5 Part C: We have to look at x = 3 and x = 5 and find the 'y' values..which is 120 and 60 (3, 120) (5, 60) Plug them into the slope formula: m=y2−y1/x2−x1 m=60−120/5−3 Simplify

oregonduck (oregonduck):

not really

oregonduck (oregonduck):

for part c is that the answer?

oregonduck (oregonduck):

ok

oregonduck (oregonduck):

can you look for it i can not find it